OpenStudy (anonymous):

how to find the exact value of sin-1(-sqrt(2)/2)

OpenStudy (anonymous):

use a unit circle

OpenStudy (anonymous):

ok I know that 7pi/4 is -sqrt(2)/2 but what does it mean by sin-1

OpenStudy (anonymous):

that usually indicates that it is looking for the inverse of the function

OpenStudy (joannablackwelder):

Yes, it is the inverse sin function.

OpenStudy (joannablackwelder):

You can write it like \[\sin ^{-1}\left( \frac{ -\sqrt{2} }{ 2 } \right)=x\]

OpenStudy (joannablackwelder):

Which is the same thing as\[\sin(x) = \frac{ -\sqrt{2} }{ 2 }\]

OpenStudy (joannablackwelder):

The inverse sin cancels out the sin to help you solve for x.

OpenStudy (joannablackwelder):

Make sense, @alyrivera ?